### v1

# Cyberattack Detection in Virtualized Microgrids Using LightGBM and Knowledge-Distilled Classifiers

Abstract
Modern microgrids depend on distributed sensing and communication interfaces, making them increasingly vulnerable to cyber–physical disturbances that threaten operational continuity and equipment safety. In this work, a complete virtual microgrid was designed and implemented in MATLAB/Simulink, integrating heterogeneous renewable sources and secondary-controller layers. A structured cyberattack framework was developed using MGLib to inject adversarial signals directly into the secondary control pathways. Multiple attack classes were emulated, including ramp, sinusoidal, additive, coordinated stealth, and denial-of-service behaviors. The virtual environment was used to generate labeled datasets under both normal and attack conditions. The datasets trained Light Gradient Boosting Machine (LightGBM) models to perform two functions: detecting the presence of an intrusion (binary) and distinguishing among attack types (multiclass). The multiclass model attained 99.72% accuracy and a 99.62% F1-score, while the binary model attained 94.8% accuracy and a 94.3% F1-score. A knowledge-distillation step reduced the size of the multiclass model, allowing faster predictions with only a small drop in performance. Real-time tests showed a processing delay of about 54–67 ms per 1000 samples, demonstrating suitability for CPU-based edge deployment in microgrid controllers. The results confirm that lightweight machine-learning–based intrusion-detection methods can provide fast, accurate, and efficient cyberattack detection without relying on complex deep-learning models. Key contributions include: (1) development of a complete MATLAB-based virtual microgrid, (2) structured attack injection at the control layer, (3) creation of multiclass labeled datasets, and (4) design of low-cost AI models suitable for practical microgrid cybersecurity.
I Introduction
Power systems are becoming more decentralized as renewable energy sources and distributed energy resources (DERs) continue to grow [32]. Microgrids now play a key role in supporting flexible, resilient, and sustainable power delivery. However, the closer connection between physical equipment and cyber layers has expanded the attack surface, increasing exposure to cyber-physical threats that can disrupt stability, reduce data accuracy, and compromise operational safety [39].
Microgrids depend on frequent measurement, communication, and control signals shared across supervisory and distributed controllers. These cyber connections create potential exposure to false data injection (FDI), denial of service (DoS), and stealth attacks [28]. Such threats may disrupt control loops, lead to incorrect state estimation, or obstruct coordination among distributed generators. Therefore, reliable intrusion detection is necessary to maintain safe real-time operation.
Traditional intrusion detection methods that use rule-based logic or signal processing often fail to provide adequate protection, particularly when facing smart or hidden attacks. Recent research has therefore shifted toward data-driven cybersecurity solutions. Machine learning based detectors show strong potential because they can learn nonlinear patterns in power data and identify anomalies with higher accuracy [3]. Progress in knowledge distillation (KD) [19] and gradient boosting models such as LightGBM [21] support the development of fast and compact classifiers for edge-level protection.
Although deep learning models such as convolutional neural networks (CNNs) and long short-term memory networks (LSTMs) can achieve high detection accuracy, they often require large computational resources. This limits their use in real-time field controllers. In contrast, lightweight non-neural models such as gradient boosted decision trees can deliver similar accuracy with faster computation and lower memory demand. These strengths make them strong candidates for intrusion detection in microgrids with limited hardware.
This work develops a complete virtual microgrid in MATLAB Simulink using cyber-physical modeling methods described in [39]. The system includes diverse renewable generation sources and secondary control layers. A structured attack injection setup was created to introduce several adversarial behaviors, including ramp, sinusoidal, additive, coordinated stealth, and denial of service attacks, into the secondary controller interface. Labeled datasets from normal and attack conditions were then exported and used to train supervised LightGBM models for binary intrusion detection and multiclass attack classification. In contrast to earlier studies such as the PSERC S82G resiliency framework [39] and the cyber-secure distributed secondary control architecture of Rath et al. [36], which mainly describe the microgrid control structure and communication vulnerabilities, this work offers a complete cyber-physical simulation and a lightweight machine learning detection system. A high-resolution virtual microgrid model with a $2~\mu\text{s}$ switching step is built and includes controlled secondary control attack scenarios. Realistic labeled datasets are created across seven operating modes, and both binary and multiclass LightGBM models are trained and can run in real-time on a regular CPU. Knowledge distillation is applied to compress the model while keeping its accuracy. This makes it suitable for use on embedded microgrid controllers. To the best of our knowledge, no earlier study brings together high-fidelity cyber-physical modeling, structured secondary control attack paths, and efficient tree boosted models for real-time intrusion detection in microgrids.
The remainder of this paper is organized as follows: Section II focuses on the review of Cyberattack Detection Techniques for Microgrids. Section III describes the virtual microgrid and attack-injection architecture. Section IV describes the dataset generation and machine learning methodology. Section V reports the evaluation results. Section VI reviews computational efficiency, and Section VII concludes the paper.
II Related Work
Research on microgrid cybersecurity has grown quickly because increased digital connectivity exposes control systems to malicious activity. Early studies showed that false data injection attacks can bypass traditional bad data detection, harming state estimation accuracy and threatening reliable operation [28]. Other work demonstrated that denial of service attacks affecting communication channels can degrade secondary control performance, causing voltage and frequency instability in inverter-based networks [29]. These findings highlighted the need for stronger cyberattack detection in cyber-physical power systems.
Early defense strategies used statistical indicators, consistency checks, and signal processing techniques to identify unusual sensor behavior [23]. These methods can handle sudden or large disturbances, but they struggle to detect coordinated or covert attacks. As a result, machine learning based intrusion detection systems have received considerable attention. For example, Sahoo and Kamwa used random forest models on real-time measurement data and reported stronger performance than conventional detection methods [41].
More recent studies have used deep learning to capture nonlinear and time-based patterns during cyber-attacks. Dey and Gupta developed a deep neural network framework that accurately detects malicious activity in smart grids [35]. Gunduz and Das used a hybrid convolutional neural network (CNN) to detect malicious behavior in smart grid metering data, achieving high attack classification accuracy while keeping false alarms low [15]. Dehghani et al. developed a deep learning intrusion detection framework for Direct Current (DC) microgrids, showing strong ability to recognize cyber-attack signatures under changing operating conditions [11]. Although these models provide strong accuracy, they require significant computational and memory resources, making real-time operation at edge controllers difficult.
To reduce complexity, lightweight learning models have been investigated. Gradient boosting methods such as LightGBM have shown strong results on structured grid data and can run well on standard Central Processing Unit (CPU) hardware [21]. XGBoost has also been used to detect abnormal events in electric networks, giving good accuracy and low response time [9]. In addition, knowledge distillation has been used in cyber-physical settings to shrink large models, allowing smaller student models to learn from larger teacher models while using less computation [19], [13]. Liu et al. showed that using knowledge distillation with federated learning (FL) can greatly reduce communication and computation demands, allowing real-time intelligence at edge cyber-physical system (CPS) nodes with limited resources [27].
Despite this progress, several limitations remain. Many current studies rely on static datasets instead of data generated from realistic microgrid simulations that include varied distributed energy resources. In addition, structured attack injection at the secondary control interface is often not implemented, limiting the evaluation of cyber-physical interactions. Finally, combining lightweight boosted models with knowledge distillation for multiclass cyberattack detection has received limited attention. These gaps motivate work that builds a MATLAB-based virtual microgrid, injects multiple types of cyberattacks at the control pathway, produces labeled datasets, and trains LightGBM classifiers enhanced through knowledge distillation for low-cost, real-time intrusion detection.
A summary of key methods is given in Table I. Although these techniques show good results with Machine Learning (ML) and Deep Learning (DL) models in different smart-grid settings, only a small number of studies have examined lightweight boosted models with knowledge distillation for multiclass cyberattack detection in virtual microgrid environments.
TABLE I: Taxonomy of Selected Smart-Grid Cyberattack Detection Approaches
Ref. Method Target Area Domain Platform
[8] ML
Disturbance &
attack detection
Power
Systems
Python
[9] SSL-DL
Cyberattack &
detection
Smart Grid Python
[10] Hybrid-CNN
Energy Theft &
detection
AMI Python
[11] DL+WSVD
DC microgrid cyber&
attack detection
DC
Microgrid
MATLAB/
Python
[12] GBDT (XGBoost)
Anomaly &
classification
Distribution Python
[14] KD+FL
Edge &
Intelligence
Alot/CPS Python
GBDT – Gradient Boosted Decision Trees, AIoT – Artificial Intelligence + Internet of Things, AMI – Advanced Metering Infrastructure, WSVD – Wavelet-based Singular Value Decomposition, SSL – Semi-Supervised Learning.
III Virtual Microgrid AND Cyberattack Injection Architecture
III-A Virtual Microgrid Overview
A detailed AC microgrid model was created in MATLAB/Simulink to represent a low-voltage distribution system with several inverter-based distributed generators (DGs). The structure follows well-known cyber–physical microgrid designs, such as the PSERC resiliency model by Srivastava et al. [39] and the cyber-secure distributed secondary-control architecture described by Rath et al [36]. As shown in Fig. 1, the model is expanded to include more DG units so that it can generate richer data for training machine-learning intrusion-detection models. The electrical system contains ten three-phase voltage-source inverters. Each inverter includes an LC output filter, feeder impedance, and standard current and voltage control loops implemented in the synchronous dq reference frame. Primary P–f and Q–V droop controllers manage power sharing among the DGs, while a distributed secondary controller restores the system’s voltage and frequency to their reference values. This layered control structure is consistent with widely used microgrid control approaches [29], [23]. The cyber layer receives measurements such as bus voltages $V_{1}$, $V_{2}$, $V_{3}$, line currents $I_{1}$, $I_{2}$, $I_{3}$, and each DG’s active power $P_{\mathrm{DGi}}$, reactive power $Q_{\mathrm{DGi}}$, and estimated frequency $f_{\mathrm{DGi}}$, for i=1,…,10. These measurements act both as inputs to the controllers and as features for the intrusion-detection dataset.
All simulations use a fixed step size of $2~\mu\text{s}$, producing 500,001 samples per scenario during a one-second time window. Each simulation starts in normal operating conditions until the microgrid reaches steady state; afterward, different cyberattacks are applied to the secondary-control channels. The resulting time-series data of voltages, currents, real and reactive power, and frequencies (V,I,P,Q,f) are saved in Comma-separated values (CSV) format. Each dataset contains the following columns:
time, $V_{1}$, $V_{2}$, $V_{3}$, $I_{1}$, $I_{2}$, $I_{3}$, $P_{\mathrm{DG1}}$, $Q_{\mathrm{DG1}}$, $f_{\mathrm{DG1}}$, …, $P_{\mathrm{DG10}}$, $Q_{\mathrm{DG10}}$, $f_{\mathrm{DG10}}$, label
The label column shows whether the system is in normal operation or experiencing a specific type of attack. These high-resolution multivariate signals form the virtual sensor dataset used to train the machine-learning models described in Section IV.
Figure 1: Architecture of the 10-DG virtual AC microgrid, showing the cyber layer, distributed secondary controller ($\xi_{i}$, $\zeta_{i}$ consensus), and local P–f / Q–V droop controllers.
III-B Primary and Secondary Control Architecture
The microgrid operates under a hierarchical control structure with two main layers, which include primary droop control at each DG and a distributed secondary controller that restores the voltage and frequency to their nominal values. This structure is based on the PSERC resiliency framework described by Srivastava et al. [39] and the cyber-secure distributed control architecture proposed by Rath et al. [36].
In this work, the architecture is extended to support ten DGs and the cyberattack scenarios considered in our study. Fig. 2 shows the step-by-step control setup for each DG, where the secondary controller adjusts the droop settings, the PLL keeps the system in sync, and the voltage and current controllers create the Pulse-width modulation (PWM) signals that control the DG’s output.
Figure 2: Cascaded control architecture of an inverter-based distributed generator (DG) unit, showing secondary control inputs, droop regulation, PLL synchronization, voltage and current control loops, and PWM inverter interface.
III-B1 Primary Droop Control
Each inverter-based DG uses standard P–f and Q–V droop control to share load without fast communication links [39]. For DG i, the droop equations are:
$\omega_{i}=\omega^{\ast}-m_{i}(P_{i}-P_{i}^{\ast}),$ (1)
$V_{i}=V^{\ast}-n_{i}(Q_{i}-Q_{i}^{\ast}),$ (2)
where $\omega^{\ast}$ and $V^{\ast}$ are the nominal frequency and voltage,
$P_{i}$ and $Q_{i}$ are the measured real and reactive power,
$P_{i}^{\ast}$ and $Q_{i}^{\ast}$ are the power setpoints,
and $m_{i}$ and $n_{i}$ are the droop coefficients.
Droop control enables proportional power sharing, but it causes small steady-state errors in voltage and frequency, which motivates the need for secondary control [39].
III-B2 Distributed Secondary Control
To eliminate the steady-state errors caused by droop control, each DG participates in a distributed consensus-based secondary controller. This approach follows the method introduced in [36]. For DG i, the frequency-correction dynamics are:
$\dot{\xi}_{i}=-k_{p}(\omega_{i}-\omega^{\ast})-c_{f}\sum_{j\in N_{i}}a_{ij}(\xi_{i}-\xi_{j}),$ (3)
$\omega_{i}^{\text{cmd}}=\omega_{i}+\xi_{i},$ (4)
and the voltage-correction dynamics are:
$\dot{\zeta}_{i}=-k_{q}(V_{i}-V^{\ast})-c_{v}\sum_{j\in N_{i}}a_{ij}(\zeta_{i}-\zeta_{j}),$ (5)
$V_{i}^{\text{cmd}}=V_{i}+\zeta_{i}.$ (6)
where, $\xi_{i}$ and $\zeta_{i}$ are integral correction states,
$N_{i}$ is the communication neighborhood of DG $i$,
$a_{ij}$ are the adjacency matrix weights,
and $k_{p}$, $k_{q}$, $c_{f}$, and $c_{v}$ are the secondary-control gains.
These equations follow the distributed controller in [36] but are scaled to the ten-DG microgrid used in this study.
III-B3 Communication Layer (Attack Target)
The secondary controller depends on low-bandwidth communication links between DGs. As described in [36], the following variables are shared over the network and are vulnerable to cyberattacks:
  • measured frequency $\omega_{i}$ and voltage $V_{i}$,
  • secondary correction terms $\xi_{i}$ and $\zeta_{i}$,
  • DG-to-DG synchronization signals,
  • feedback messages exchanged with the microgrid central controller (MGCC).
These communication channels form the main attack surface for the scenarios described later in Section 4.
III-B4 Inner Voltage and Current Loops
Below the droop and secondary layers, each DG uses PI-based controllers in the dq-frame [39]. The voltage loop is:
$v_{d}^{\ast}=V_{i}^{\text{cmd}}+K_{pv}(V_{d}-V_{d}^{\ast})+K_{iv}\int(V_{d}-V_{d}^{\ast}),dt,$ (7)
and the current loop is:
$i_{d}^{\text{cmd}}=I_{d}^{\ast}+K_{pc}(i_{d}-i_{d}^{\ast})+K_{ic}\int(i_{d}-i_{d}^{\ast}),dt.$ (8)
These inner loops are not directly attacked, but convert corrupted secondary-control signals into physical disturbances that appear in voltage, current, and power waveforms.
III-C Secondary Control Interface for Attack Injection
The secondary control layer is the most communication-dependent component of the microgrid architecture, making it a natural target for cyberattacks. In this work, all malicious signals are injected directly into the secondary control input channels, specifically the distributed consensus signals $\xi_{i}$ and $\zeta_{i}$ that restore frequency and voltage across all distributed generators.
  • Attacks on $\xi_{i}$ primarily distort active power and frequency regulation,
  • Attacks on $\zeta_{i}$ impact reactive power and voltage regulation,
  • Attacks on both signals emulate coordinated disturbances affecting overall stability.
Figure 3: Communication Pathways of the Secondary Control Layer With the Injection Point of the Corrupted Consensus Signals.
Fig. 3 shows the ring-based secondary-control communication network used by the ten inverter-based DG units. Each DG exchanges the consensus signals $\xi$ and $\zeta$ with its neighbors, and the attack is injected by corrupting the incoming signal received by DG1. This design choice follows the cyber-physical vulnerability analysis presented by Srivastava et al. [39] and the threat model from the cyber-secure microgrid architecture paper used in this work. Because the DGs share information only with their neighbors over a sparse communication network, an attacker who breaks into just one communication link can affect several DGs at once. Earlier research has shown that even small false signals added to the secondary controller can spread through the network, change how power is shared, and weaken the microgrid’s dynamic response [20, 3].
To study these weaknesses under realistic conditions, this work applies six types of cyberattacks. Each attack is injected into the secondary-control input of selected DGs. Fig. 4 provides a compact illustration of the six secondary-control attack profiles considered, highlighting the characteristic shape of each injected disturbance.
Figure 4: Overview of the six secondary-control cyberattack patterns applied to the consensus signal $\xi_{i}(t)$ and $\zeta_{i}(t)$.
III-C1 Ramp Attack
A gradually increasing bias is added to the distributed secondary-control frequency-correction signal $\xi_{i}$. The malicious modification is expressed as:
$\xi_{i}^{\text{mal}}(t)=\xi_{i}(t)+rt$ (9)
where r is the ramp slope. Because the frequency loop is very sensitive to slow long-term drift, even small errors can gradually build up and create an active-power imbalance. By imitating a slowly increasing load, the attacker causes the controller to shift power incorrectly. Similar slow-moving biases have been shown to destabilize secondary control loops in microgrids when not corrected [3, 27, 28].
III-C2 Slow Ramp Attack
A very small slope is injected into the secondary voltage-restoration signal. The attack modification is given by:
$\zeta_{i}^{\text{mal}}(t)=\zeta_{i}(t)+r_{s}t,\quad r_{s}\ll r.$ (10)
This attack targets $\zeta_{i}$ because the voltage loop can conceal very slow deviations, allowing the attacker to create a stealthy reactive-power imbalance without triggering threshold-based detectors. Such slowly drifting anomalies have been reported in voltage stability assessments and cyber-physical intrusion analyses [23, 3].
III-C3 Sinusoidal Attack
A coordinated oscillatory signal is injected into both secondary control inputs. The malicious change is given by;
$\xi_{i}^{\text{mal}}(t)$ $=\xi_{i}(t)+A\sin(\omega_{a}t),$ (11)
$\zeta_{i}^{\text{mal}}(t)$ $=\zeta_{i}(t)+A\sin(\omega_{a}t).$ (12)
Oscillatory disturbances affect both P and Q regulation, exciting inverter dynamics and distorting transient responses. Injecting them simultaneously into $\xi_{i}$ and $\zeta_{i}$ emulates a coherent, dynamic intrusion, similar to oscillatory false-data injections shown to degrade microgrid performance [11].
III-C4 Additive Bias Attack
A constant malicious offset is added to both secondary signal, which is given by:
$\xi_{i}^{\text{mal}}(t)$ $=\xi_{i}(t)+b,$ (13)
$\zeta_{i}^{\text{mal}}(t)$ $=\zeta_{i}(t)+b.$ (14)
This attack shifts the internal reference points of the controller, producing sustained misregulation in both active- and reactive-power sharing. Fixed-bias spoofing is a classical false-data injection approach and is rigorously analyzed in the foundational work of the literature [28].
III-C5 Coordinated Stealth Attack
In this mode, small, correlated deviations are injected across multiple DGs.
$\xi_{i}^{\text{mal}}(t)$ $=\xi_{i}(t)+f_{s}(t),$ (15)
$\zeta_{i}^{\text{mal}}(t)$ $=\zeta_{i}(t)+\alpha f_{s}(t).$ (16)
where $f_{s}(t)$ is a low-amplitude waveform designed to remain within normal envelopes.
This mode targets both control loops because coordinated multi-variable manipulation is one of the most difficult patterns to detect. By maintaining correlations between the disturbances, the attack closely mimics normal operation while degrading power-sharing performance. Similar multi-signal stealth strategies appear in recent intrusion-detection literature[35, 15].
III-C6 Denial-of-Service (DoS) Attack
A DoS attack freezes the frequency-restoration signal by preventing new consensus data from reaching DGs. This malicious modification is given by:
$\xi_{i}^{\text{mal}}(t)=\xi_{i}(t_{k}),\qquad t>t_{k}.$ (17)
The frequency channel is targeted because loss of synchronization typically manifests first through degraded P–f restoration. With stale control data, DGs drift apart in frequency and voltage, leading to instability. DoS problems in microgrid control are well documented in the literature, with similar observations reported in broader cybersecurity surveys [3, 29].
All attacks were injected exclusively into the secondary-control correction layer, ensuring that disturbances propagate through the cyber-control channel rather than through sensors or inverter hardware. This modeling approach aligns with real microgrid vulnerabilities documented in recent literature [39, 3, 29]. All simulations were run in MATLAB/Simulink using MGLib within the custom virtual microgrid built for this study.
III-D MATLAB/Simulink Implementation and Data Logging Framework
The virtual microgrid and cyberattack-injection environment were implemented entirely in MATLAB/Simulink using custom Simscape-based inverter models, hierarchical controllers, and structured cyberattack subsystems. Two simulation environments, Normal.slx and MGLib.slx, were used. Normal.slx implements the full 10-DG microgrid and executes all cyberattack scenarios, while MGLib.slx provides a library of reusable distributed-generation and control modules used to construct and parameterize the virtual microgrid. The modeling philosophy follows the cyber-physical microgrid resilience principles and the distributed secondary control architecture outlined in the literature [39, 28].
III-D1 Model Structure
As illustrated in Fig. 1, the Simulink model follows the physical, control, and cyber–physical behavior of an inverter-dominated microgrid and is organized into four connected layers.
a) Physical Layer
The physical layer contains all inverter-based DG units, along with the AC distribution lines and system loads. Each DG uses a voltage-source inverter model with droop-based primary control, following the standard methodology presented in [29]. The switching behavior is fully retained, allowing realistic harmonic and dynamic responses.
b) Secondary Control Layer
A distributed consensus-based voltage and frequency restoration controller regulates the network. Each DG exchanges correction signals $\xi_{i}$ and $\zeta_{i}$ with its neighbors. The controller setup is based on the cyber-secure distributed control framework introduced in [36], which describes how communication links, coordination graphs, and vulnerability paths operate in autonomous AC microgrids.
c) Cyber Layer (Communication Graph)
Peer-to-peer communication links between DGs are implemented using Simulink signal routing and discrete update blocks. Communication irregularities such as delay, sample jitter, and packet manipulation are introduced at this layer. These imperfections reflect realistic cyber-physical interactions, consistent with observations from recent microgrid cybersecurity reviews [1].
d) Attack-Injection Layer
This layer contains structured cyberattack blocks that inject ramps, biases, oscillations, or packet-freeze events directly into the control corrections $\xi_{i}$ and $\zeta_{i}$. The attack signals follow the formulations defined in Section III-C and are designed to propagate through the entire secondary-control loop in a physically meaningful manner.
III-D2 Attack Injection Blocks
Each of the six cyberattacks defined in Section III-B was implemented as an independent Simulink subsystem so that scenarios can be executed in a controlled and repeatable way. The subsystems realize the ramp, slow-ramp, sinusoidal, additive-bias, coordinated stealth, and DoS profiles using standard integrator, sine-wave, offset, and sample-and-hold blocks following the formulations in (9)–(17).
This set of attacks follows the cyber-physical disturbance pathways described in [39] and the microgrid cyberattack case studies summarized by [3].
III-D3 High-Frequency Simulation Parameters
All simulations were executed using high-fidelity switching resolution:
  • Sample time: $2~\mu\text{s}$
  • Solver: Fixed-step, discrete
  • Switching model: High-resolution PWM behavior
  • Simulation duration: 1.0 s
  • Step size: $2\times 10^{-6}\text{ s}$
Such small time steps follow best practices for simulating inverter-dominated microgrids. They allow the model to capture switching ripple and fast system dynamics accurately, which agrees with commonly accepted transient-modeling guidelines [24, 16].
III-D4 Data Logging and Export Pipeline
A dedicated logging subsystem recorded all cyber-physical signals from every DG:
  • Voltage $V_{i}$, Current $I_{i}$, Real/reactive power $P_{\mathrm{DG1}}$, $Q_{\mathrm{DG1}}$, Frequency estimate $f_{\mathrm{DG1}}$
  • Attack state (binary label: 0 = normal, 1 = attack)
  • Attack type index (Normal, Ramp, Slow Ramp, Additive, Sinusoid, Stealth, DoS)
  • Time index
The signals were saved using To Workspace and To File blocks in array form, and then exported to CSV files using MATLAB’s built-in data-export functions.
Each simulation produced about 500,001 samples across 38 variables, resulting in high-resolution datasets that are suitable for machine-learning analysis.
III-D5 Dataset Consolidation
All scenario outputs (Normal and six attack types) were exported as CSV files. These files were later merged and prepared into a unified supervised-learning dataset, as detailed in Section IV.
III-E Noise Representation and Robustness Modeling
Although cyberattacks introduce significant disturbances into the microgrid secondary control layer, inverter-based microgrids naturally operate under several sources of physical and numerical noise even in the absence of malicious behavior. Because these disturbances influence voltage, current, power, and frequency measurements, they directly affect the data used to train machine learning models. To ensure that the learning algorithms developed in this work generalize to real-world operating conditions, the virtual microgrid was simulated using a switching-level model with a fixed $2~\mu\text{s}$ step size. As depicted in Fig. 5, this configuration inherently produces structured noise and variability consistent with practical microgrid deployments, consistent with the recommendations in [39] and the cyber-physical control behavior described in the reference microgrid architecture.
Figure 5: Illustration of noise and discretization effects in the virtual microgrid, showing PWM switching ripple in voltage and current signals, quantization steps from digital measurement, and small timing jitter in the consensus updates of the secondary controller.
III-E1 Measurement Noise from High-Fidelity Switching Dynamics
Each distributed generator uses a PWM-based voltage source inverter modeled at microsecond resolution. As noted in prior transient modeling studies of power-electronic systems [7], switching ripple introduces small but persistent harmonic components into voltage and current waveforms [16]. These distortions act as a realistic source of sensor noise and ensure that the resulting dataset reflects the non-ideal characteristics of real converter-dominated microgrids. Unlike artificially smoothed or phasor-domain models, this switching representation prevents the classifier from overfitting to unrealistically clean signals.
III-E2 Numerical Discretization Noise in Sensing and Control
The microgrid uses sample-based measurement of real power, reactive power, and frequency at each DG. Because these quantities are computed from discrete-time signals, they exhibit quantization effects and numerical jitter [43]. Such effects mimic the finite precision and sampling behavior of real-world measurement systems and digital controllers. This type of noise is particularly important when classifying subtle or slowly varying attacks such as Slow Ramp and Coordinated Stealth disturbances, which must be detected against a background of small natural fluctuations [39].
III-E3 Communication Jitter in Distributed Secondary Control
The distributed secondary control layer operates through discrete communication exchanges among DG controllers. Although no explicit random delay blocks were added, the discrete-time update behavior of Simulink inherently produces small timing mismatches between controller nodes. As described in [39] and prior reviews of microgrid cyber-physical vulnerabilities [3], such jitter is representative of real-world low-bandwidth microgrid communication networks. This adds realistic uncertainty to consensus variables $\xi_{i}$ and $\zeta_{i}$, improving the resilience of the classifier to benign fluctuations in the absence of attacks.
III-E4 Implications for Machine Learning Robustness
The combination of switching ripple, discretization noise, and communication jitter forms a realistic disturbance environment. These factors help ensure that the LightGBM models trained in later sections do not overfit to clean or idealized trajectories. Instead, they learn to classify attacks under noisy, non-linear, and physically accurate operating conditions, consistent with recommendations in prior microgrid intrusion detection studies [3, 11, 12]. This contributes directly to the generalization ability shown in Section V, where the classifier achieves high accuracy across all attack modes.
IV Dataset Generation and Machine Learning Methodology
This section describes how high-resolution microgrid simulation data were converted into structured datasets for supervised learning, and how the machine-learning models were designed, trained, and optimized. All data originate from the MATLAB/Simulink virtual microgrid developed in Section III, where cyberattacks were injected directly into the distributed secondary-control layer.
IV-A High-Resolution Data Generation in Simulink
Each scenario, whether normal or under one of the six attack modes, was simulated for 1 s using a $2~\mu\text{s}$ fixed-step solver, yielding 500,001 time samples. Small time steps are recommended for inverter-dominated microgrids because they preserve fast switching behavior, harmonic content, and non-linear transient dynamics that influence cyber-physical stability, as emphasized in modern power-electronics transient modeling studies [37, 34]. For each time step, the following measurements were logged:
  • Three-phase bus voltages: $V_{1}(t)$, $V_{2}(t)$, $V_{3}(t)$
  • Three-phase line currents: $I_{1}(t)$, $I_{2}(t)$, $I_{3}(t)$
  • Active and reactive power of $DG_{i}$, $i\in\{1,\dots,10\}$: $P_{\mathrm{DG}i}(t),\;Q_{\mathrm{DG}i}(t)$
  • Local frequency measurements: $f_{\mathrm{DGi}}(t)$
These variables form a virtual sensor layer analogous to what would be measured by Phasor measurement units (PMUs), smart inverters, or microgrid controllers in practice.
Each simulation also included:
  • Attack state (binary label): $y_{\text{bin}}\in\{0~\text{(normal)},\,1~\text{(attack)}\}$
  • Attack type index (Normal, Ramp, Slow Ramp, Additive, Sinusoid, Stealth, DoS)
IV-B Dataset Consolidation and Attack Annotation
After simulation, all CSV files were merged into a single dataset. Scenario filenames were parsed to automatically assign:
  • Binary Label:
    $y_{\text{bin}}=\begin{cases}0,&\text{normal}\\ 1,&\text{attack}\end{cases}$
  • Multiclass Attack Type:
    $y_{\text{multi}}\in\{0:\text{Normal},\;1:\text{Additive},\;2:\text{Ramp},\;3:\text{Slow Ramp},\;4:\text{Sinusoid},\;5:\text{Stealth},\;6:\text{DoS}\}$
  • The combined dataset has a size of $500{,}001\times 7\approx 3.5$ million samples.
    This included 38 continuous features and two target labels.
IV-C Memory-Aware Downsampling Strategy
Training on the full 3.5-million–sample dataset exceeded available GPU/CPU memory. To retain fidelity while reducing size, a balanced downsampling scheme was used [8]. All attack samples were kept, and only 10–-15% of normal samples were retained. This achieved a 50-–60% reduction while preserving transients at attack onset, normal–attack boundaries, and temporal dynamics essential for classification. The final working dataset contained approximately 1,030,197 samples.
IV-D Noise-Resilient Feature Normalization
Because the microgrid includes switching ripple, discretization noise, and communication jitter (Section III), normalization must preserve subtle variations. To avoid memory overflow, incremental (chunk-based) normalization was adopted using the standard z-score formula [26]:
$x_{\text{norm}}=\frac{x-\mu}{\sigma}$ (18)
where $x$ denotes the raw feature value, $\mu$ is the feature mean, and $\sigma$ is the corresponding standard deviation.
The key points to note are that the Data were converted to float32 to reduce the memory footprint. Large numerical fields were rescaled (e.g., division by 1000) to avoid overflow, and normalization was computed in chunks to avoid Random Access Memory (RAM) exhaustion. The resulting dataset had $\mu\approx 0\;\text{and}\;\sigma\approx 1$ for all features.
IV-E Train–Validation–Test Split
A stratified split was used to keep all attack types represented in the same proportions across the training, validation, and test sets [6]. The training set contains 721,137 samples, while the validation and test sets each contain 154,530 samples. Stratification covered both Binary label and Multiclass attack type to avoid bias and maintain class balance.
IV-F Binary Attack Detection Using LightGBM
Binary classification separates normal operation from attack conditions. LightGBM is used because it is fast on structured data, requires very little memory, performs extremely well without Graphics processing unit (GPU) acceleration, and has a strong track record in cyber-physical system applications [21]. The model minimized the standard binary log-loss:
[…]
where $y\in\{0,1\}$ is the ground–truth binary label (0 = normal, 1 = attack) and $\hat{p}$ is the predicted probability that the sample belongs to the attack class.
IV-G Multiclass Attack-Type Classification
A second LightGBM model was trained to classify all 7 attack types at the same time. The goal was for the model to correctly tell the difference between each type of cyberattack. To achieve this, the model minimized the standard multiclass cross-entropy (log-loss) function:
$L_{\text{mc}}=-\sum_{k=1}^{7}y_{k}\ln\left(\hat{p}_{k}\right)$ (20)
where $y_{k}$ is the one–hot encoded ground–truth label and $\hat{p}_{\,k}$ is the predicted probability for class $k$. The confusion matrix (Section V) confirms nearly perfect class separability.
IV-H Knowledge Distillation for Lightweight Deployment
To reduce inference time and model size, knowledge distillation (KD) was used [19].
A teacher–student architecture was implemented:
  • Teacher: full LightGBM multiclass model
  • Student: smaller LightGBM model trained using Hard labels, Soft probabilities from the teacher, and Temperature-scaled softmax.
$L=\alpha\,CE(y_{\text{hard}},s)+\beta\,KL\!\left(\mathrm{softmax}(T_{t}),\,\mathrm{softmax}(T_{s})\right)$
where:
  • $t$ and $s$ are the teacher and student logits,
  • $T$ is the distillation temperature,
  • CECE is the standard cross-entropy loss using hard labels,
  • KLKL is the Kullback–Leibler divergence between teacher and student softened probabilities,
  • $\alpha$ and $\beta$ control the weight of the hard-label and soft-label losses.
The LightGBM-based architecture meets real-time requirements and achieves efficiency gains consistent with findings in [21, 13].
V Results
This section evaluates the proposed cyberattack-detection framework, covering binary detection, multiclass attack classification, feature-importance analysis, knowledge-distillation results, and real-time prediction performance. All experiments were performed on the normalized, downsampled dataset produced from the high-resolution simulations described in Section IV.
V-A Binary Attack Detection Performance
Using the label definitions $y_{\text{bin}}$ from Section IV-B, the binary LightGBM classifier was trained on the merged dataset to discriminate between normal and malicious operating modes. After training with early stopping, the binary model achieved the results shown in tableII. This shows that the model can dependably identify whether the microgrid is operating normally or under attack, even when exposed to a wide range of disturbance scenarios.
Figure 6: Binary Confusion Matrix for Cyberattack Detection in the Microgrid
Confusion Matrix Analysis
From Fig. 6, the binary confusion matrix demonstrates:
  • Very low false-negative rates, meaning almost all attacks were detected.
  • Most errors correspond to false positives, where normal operating conditions are occasionally classified as attacks.
  • Attack samples are identified correctly in more than 99% of cases, showing that the model stays reliable even when noise and inverter switching ripple are present.
Overall, the binary classifier works well as a strong first line of defense. It can quickly detect when a cyberattack is happening with very high sensitivity.
V-B Multiclass Attack-Type Classification
To identify specific attack types, a multiclass LightGBM model was trained using the label definitions $y_{\text{multi}}$ from Section IV-B.
V-B1 Test-Set Performance
These results of the multiclass classifier, as summarized in tableII, indicate almost perfect classification across all attack types.
TABLE II: Performance Comparison of Binary, Multiclass, and Distilled Models
Model Accuracy (%)
Macro
F1 (%)
Weighted
F1 (%)
Inference
(ms/1000)
Binary
LightGBM
94.8 94.3 94.3 53.96
Multiclass
LightGBM
99.72 99.62 99.72 67.00
Student (KD) 99.70 99.70 99.70 $\sim$18.0
V-B2 Confusion Matrix Interpretation
From Fig. 7, the multiclass confusion matrix shows:
Figure 7: Multiclass Confusion Matrix for Cyberattack Detection in the Microgrid
  • Strong diagonal dominance, meaning the model is very good at telling each attack type apart.
  • Very few mistakes, even between Ramp and Slow-Ramp attacks, which normally look similar.
  • Perfect or near-perfect recognition of Additive, Sinusoid, Stealth, and DoS attacks.
The strong separation between attack types shows that the dataset provides enough unique signals for each mode. This also shows that the model can clearly recognize these differences in the data and use them to classify each attack type accurately.
V-C LightGBM Hyperparameters
The binary and multiclass classifiers were implemented using LightGBM. All model configurations, including the boosting setup, learning parameters, and iteration counts, are summarized in Table III. The binary model employed a standard binary objective, while the multiclass model used a seven-class formulation with early stopping determining the final number of boosting rounds. A compact student model was also trained using knowledge distillation. As shown in Table III, the student configuration uses reduced tree complexity and fewer boosting rounds to enable lower memory usage and faster inference while maintaining strong predictive performance.
TABLE III: LightGBM Hyperparameters for Binary, Multiclass, and Student Models
Parameter Binary Multiclass Student
Objective binary multiclass multiclass
Number of classes 7 7
Boosting type gbdt gbdt gbdt [^1] Parameters not explicitly set for the student model use LightGBM default values.
Num. leaves 63 63 15
Learning rate 0.05 0.05 0.10
Feature fraction 0.9 0.9 0.8
Bagging fraction 0.8 0.8 0.8
Bagging frequency 5 5  [^1] Parameters not explicitly set for the student model use LightGBM default values.
Num. iterations (set) 200 200 50
Best iteration 200 104 50
Threads (n_jobs) 1-1 1-1  [^1] Parameters not explicitly set for the student model use LightGBM default values.
V-D Statistical Robustness and Ablations
Robustness was assessed using per-class F1-scores and targeted feature-removal tests. The Normal mode remains the most difficult to classify (F1 $\approx$ 0.983), while all attack classes exceed 0.996, indicating stable performance even under subtle secondary-control disturbances. As summarized in Table IV, removing active-power, reactive-power, or frequency signals produces negligible changes in macro-F1 (0–0.01% drop), showing that these quantities are highly redundant across distributed generators. In contrast, removing only the three bus-voltage measurements (V1–V3) decreases macro-F1 by approximately 1 percentage point, identifying voltage as the most informative feature group for distinguishing normal and ramp-type behaviors. Line-current removal shows no measurable effect. Overall, the classifier remains highly robust, with voltage magnitudes being the only feature group that provides uniquely discriminative information for this attack set.
TABLE IV: Ablation Study: Effect of Removing Each Feature Group
Feature Removed F1 Drop Interpretation
None (Baseline) 0.9948 Full feature set.
$P_{\mathrm{DG}}$ (Active Power) 0.9948 0.00% No measurable impact.
$Q_{\mathrm{DG}}$ (Reactive Power) 0.9948 0.00% Redundant information.
$f_{\mathrm{DG}}$ (Frequency) 0.9947 0.01% Minor contribution.
$V_{1}$$V_{3}$ (Bus Voltage) 0.9847 1.01% Most informative group.
$I_{1}$$I_{3}$ (Line Currents) 0.9948 0.00% Redundant with others.
V-E Feature-Importance Analysis
Gain-based feature importance was computed for the multiclass classifier to understand which physical variables contribute most to attack detection.
Key observations include:
  • Active power measurements $P_{\mathrm{DG}i}$ were dominant predictors, which aligns with the fact that many attacks influence power-sharing dynamics.
  • DG local frequencies $F_{\mathrm{DG}i}$ were critical for detecting oscillatory disturbances (Sinusoid) and synchronization-affecting attacks (Stealth, DoS).
  • Line currents $I_{1}$, $I_{2}$, $I_{3}$ showed high importance for detecting abrupt disturbances such as DoS and Ramp.
  • Voltage measurements played a consistent but moderate role across all classes.
Interpretation
The feature-importance distribution indicates that the classifier relies on microgrid-wide behavioral signatures, not isolated measurements. Because it depends on many variables, it is more resilient and less likely to be misled by an attacker manipulating a single sensor.
V-F Knowledge Distillation Evaluation
KD process was used to compress the multiclass LightGBM model into a lighter student model suitable for real-time embedded controllers [19].
Performance Comparison
As shown in Fig. 8, the student achieved similar results as the teacher model.
Figure 8: Teacher–student model comparison for microgrid cyberattack detection.
The knowledge-distilled student preserves diagnostic performance while reducing inference time by approximately 73% relative to the teacher model. This makes the detector suitable for microgrid edge controllers with limited processing power.
Implication
This demonstrates that knowledge distillation is an effective strategy for deploying cyberattack detection on microgrid controllers with limited computational resources.
V-G Real-Time Prediction Behavior
To assess the real-time decision capability of the proposed intrusion-detection model, a set of ten randomly selected unseen samples was extracted from the held-out test dataset. Because the test set was generated directly from the labeled Simulink scenarios described in Section IV, each sample retains an associated ground-truth attack label that reflects the exact attack mode injected during simulation. These labels were not exposed to the model during training or validation.
Figure 9: Real-time multiclass attack prediction on unseen test samples.
As shown in Fig. 9, the trained multiclass LightGBM model was then used to make a prediction for each selected sample, one point at a time. Each predicted label was directly compared with its corresponding ground-truth label in the test set, and all ten predictions matched perfectly. This confirms that the classifier generalizes to unseen temporal slices and decision boundaries remain stable under noise and switching ripple. The system behaves reliably for point-wise real-time inference.
VI Computational Efficiency and Deployment Considerations
Effective cyberattack detection in microgrids must be both accurate and fast [22]. The detection algorithm also needs to run with low computational cost and meet strict real-time requirements [31]. This section evaluates the computational efficiency of the LightGBM-based detection framework, including model size, inference speed, CPU-only performance, and the improvements gained through knowledge distillation. All runtime measurements correspond to the same experimental setup used in Section V.
VI-A Hardware Environment
All inference-time and model-size measurements were performed on a virtualized Google Colab CPU instance. The system uses a GenuineIntel Intel Xeon @ 2.20 GHz processor with the following configuration:
  • 1 physical core, 2 threads
  • x86_64 architecture
  • KVM hypervisor virtualization
  • 55 MB L3 cache
  • AVX, AVX2, FMA extensions enabled
This environment represents a modest CPU-only computing platform, allowing fair comparison with embedded-class processors that lack GPU acceleration.
VI-B Model Size and Memory Footprint
All models, including the binary classifier, multiclass classifier, and distilled student model, were exported in the standard LightGBM text format for size evaluation. Because LightGBM represents each tree using simple threshold-based splits rather than dense numerical tensors, its memory footprint is inherently small compared with deep learning models [21].
Key observations include:
  • Binary LightGBM model: smallest size because it learns only a 2-class boundary.
  • Multiclass LightGBM model: 4.83 MB, reflecting the larger number of classes and trees.
  • Distilled student model: 0.60 MB, an 87.7% reduction, achieved by using fewer leaves and shallower trees while preserving accuracy.
These models are much smaller than common deep-learning intrusion detectors, like CNN or LSTM models used for grid event detection, which often range from a few megabytes to hundreds of megabytes [4].
Practical implication:
All three LightGBM models fit easily within the memory limits of low-power embedded hardware, including ARM Cortex-M microcontrollers, TI DSP control boards, Industrial IoT processors and Edge devices like Raspberry Pi units. This small memory size allows the intrusion detection system to run directly on microgrid controllers, without needing cloud support or GPU hardware [5, 2, 33].
VI-C Inference-Time Benchmarking
Inference latency was evaluated using 1000 consecutive unseen samples from the test dataset. Table IV summarizes the measured times. The binary model is fastest because it performs a single 2-class decision, while the multiclass model is slightly slower due to additional tree evaluations. The student model is the fastest overall, confirming the computational advantage of knowledge distillation.
Real-Time Feasibility
Microgrid control loops typically operate with 100–200 ms update intervals for secondary controllers [14] and 20–50 ms recommended detection window for cyberattack monitoring [30]. All models meet these real-time constraints comfortably, even on CPU-only hardware, enabling fully embedded deployment without GPUs.
VI-D 3 CPU-Only Deployment
A major benefit of the proposed framework is its compatibility with CPU-only environments. Unlike CNNs, LSTMs, or transformer-based detectors that need many floating-point operations and often rely on GPU acceleration, LightGBM uses simple operations such as threshold comparisons, integer additions, and shallow tree traversal. These operations are extremely efficient on embedded processors.
In contrast, CNN-based smart grid detectors [18, 42] rely on millions of convolution operations. LSTM-based microgrid models 11 require large matrix multiplications and memory-intensive time-state storage. The proposed LightGBM+KD framework avoids these bottlenecks entirely.
VI-E Knowledge Distillation for Edge Intelligence
KD was incorporated to reduce computational load while preserving classification fidelity. A smaller student LightGBM model was trained to mimic a larger teacher model using softened probability targets and temperature scaling. The results from Section V align with findings in recent edge-intelligence and resource-efficient ML studies [10, 44] confirming that KD lowers RAM usage, reduces energy consumption, improves inference speed, increases resilience to noisy measurements, and enables real-time deployment in cyber-physical systems. Thus, KD enhances the practicality of the proposed framework for real-world microgrid controllers.
VI-F Scalability Considerations
The architecture exhibits strong scalability due to the properties of tree-based models and the modularity of the virtual microgrid pipeline.
Scalability Drivers
VI-F1 Inference does not rely on differential equations
Unlike physics-based estimators, LightGBM classifiers require no real-time numerical integration.
VI-F2 Tree-based models scale sublinearly
Adding new DG measurements or new power-quality features only increases the model size modestly [21].
VI-F3 Distributed deployment
Identical local detectors can be placed into each DG controller [17].
VI-F4 Extendable taxonomy
New cyberattack modes can be integrated simply by retraining with additional labeled datasets.
VI-F5 Compatibility with federated learning
Techniques similar to [25] can be applied so that decentralized microgrids collaboratively improve the detection model without sharing raw data.
VI-G Implication for Future Systems
The results demonstrate that the proposed detection framework is not limited to the specific microgrid modeled in this study. Instead, it provides a scalable foundation that can extend to a broad class of modern cyber-physical power systems. Because the LightGBM-based detectors operate with low computational overhead, require no GPU acceleration, and retain high accuracy under switching noise and communication disturbances, the approach is suitable for deployment in several emerging grid architectures [38].
Firstly, multi-DER distribution grids, which increasingly rely on inverter-interfaced resources, can benefit from having lightweight anomaly detectors co-located with local controllers [30]. The capability to run efficiently on CPU-only hardware makes on-device inference feasible even at high sampling rates. Furthermore, inverter-dominated AC and DC microgrids can integrate the framework at the secondary-control or tertiary-control layers without modifying existing control hardware [14]. This positions the model as a practical cybersecurity safeguard to accompany power-sharing and voltage-restoration mechanisms. Thirdly, the method aligns naturally with edge-to-cloud architectures. A local LightGBM model can run at the edge (e.g., on DG controllers or protection relays), while aggregated attack indicators or distilled student models may be shared with supervisory systems or cloud-hosted analytics platforms [40].
Finally, the framework is compatible with hybrid AC/DC microgrid environments, which require fast decision-making across heterogeneous converter topologies. Overall, the combination of high accuracy, low inference latency, and modest computational requirements positions the proposed system as a strong candidate for next-generation resilient smart grids. It supports real-time cyberattack detection across many types of microgrid setups and can easily scale to new distributed energy systems.
VII Conclusion
This work presented a complete cyber-physical framework for generating, labeling, and classifying secondary-control cyberattacks in inverter-based microgrids. A high-fidelity virtual microgrid was constructed in MATLAB/Simulink using distributed droop-controlled DGs, consensus-based secondary controllers, and structured cyberattack injection pathways. The simulation environment recorded detailed physical and cyber behavior at a very high time resolution of $2~\mu\text{s}$. This made it possible to generate realistic and well-labeled datasets that accurately represent both normal operation and different types of cyberattacks.
A complete machine learning pipeline was built to process the generated datasets. After applying incremental normalization and balanced downsampling, we trained binary and multiclass LightGBM models to detect and classify six representative secondary-control attack types. The multiclass model reached 99.7% accuracy, showing that the attack types can be recognized correctly by the classifier. Knowledge distillation was then used to compress the model, creating a smaller student classifier that kept the same accuracy while reducing the inference time by about 73%.
The results show that the tree-boosted models provide a strong mix of accuracy, reliability, and fast computation. Their compact structure enables practical real-time deployment across a wide range of microgrid controllers. Furthermore, the structured dataset and modeling approach provide a repeatable blueprint for future cyber-resilient microgrid studies. Future work will extend the framework to include (i) additional DER types and converter topologies, (ii) communication-layer adversaries with time-varying topologies, (iii) reinforcement-learning-based adaptive controllers, and (iv) federated or privacy-preserving architectures to enable secure multi-microgrid coordination.
Overall, this study shows that using high-resolution cyber-physical simulations together with lightweight gradient-boosting models and knowledge distillation provides a practical and effective way to detect cyberattacks in real time within modern microgrids.
Acknowledgments
Special thanks to Professor Suman Rath for his guidance and mentorship during the AI in Power System course, as well as for his continued supervision and support throughout the doctoral research project.